OpenStudy (anonymous):

find the equation of tangent line indicated point: y= square root of (3x-1 / 2x+1) at (-1,2) im doing it right now.. am i on the right track? please help

OpenStudy (anonymous):

i know i have to use the\[y-y _{1} = m(x-x _{1})\] at some point

OpenStudy (anonymous):

and the \[f(x)= \lim_{h \rightarrow 0} \frac{ f(x-h) - f(x) }{h }\] too

OpenStudy (anonymous):

i just dont know when

OpenStudy (anonymous):

so far i did